1.If a wheel is turning at 7.7rad/s, the time (in seconds) it takes to complete one revolution is about:

Select one:

a. 0.82

b. 2.5

c. 48

d. 24

e. 0.41

2.A flat (unbanked) curve on a highway has a radius of 222.8m. A car rounds the curve at a speed of 27.6m/s. What is the minimum coefficient of friction that will prevent sliding?

3.Suppose a stationary 52.5-kg box is on top of your table. What is the work done by gravity?

Select one:

a. 52.50 J

b. 0 J

c. 9.8 J

d. 514.50 J


1
Expert's answer
2022-01-07T09:28:54-0500

1. "T=2\pi/\omega=0.82\text{ s}."

2. "\mu=v^2/(gR)=0.35."

3. "W=0."